OpenStudy (yanasidlinskiy):

1. Solve the absolute value equation. Check your solutions. Can somebody check my work if I got this right?? | 5x + 13| = –7 -13 -13 5x = - 20 x = -4

OpenStudy (texaschic101):

| 5x + 13 | = -7 5x + 13 = -7 AND -(5x + 13) = -7 5x = -7 - 13 -5x - 13 = -7 5x = -20 -5x = -7 + 13 x = -4 -5x = 6 x = -6/5 I am not sure about this....you need to get a second opinion

OpenStudy (anonymous):

Stupid Moderators, they just gave me a warning for what you just did.

OpenStudy (texaschic101):

crazy, huh....I got suspended before for doing this.

OpenStudy (texaschic101):

the first solution...-4 works, but the second one -6/5 does not. So I do not know if the answer is no solution or not ??

OpenStudy (yanasidlinskiy):

Its ok. i ll find the answer somewhere else:)

OpenStudy (texaschic101):

sorry :(

OpenStudy (yanasidlinskiy):

But thanks for your help though:)